When 3.949 grams of a hydrocarbon, CxHy, were burned in a combustion analysis apparatus, 12.39 grams of CO2 and 5.073 grams of H2O were produced.

In a separate experiment, the molar mass of the compound was found to be 56.11g/mol. Determine the empirical formula and the molecular formula of the hydrocarbon.

  • Enter the elements in the order presented in the question.
    empirical formula =
    molecular formula =
Answer #1

Convert moles to grams by dividing by the Molecular weight:

Moles CO2=12.39/44.01=0.2815

Moles H2O=5.073/18.01=0.0.2815

Moles C=0.2815

Moles H = 0.2815×2= 0.563

Ratio C:H by moles = 0.2815 : 0.563⇒

1:2

So the empirical formula is CH2

Molecular weight of the compound is 56.11g/mol

Empirical formula molecular weight = 14

Thus Molecular formula = empirical formula * 4

= CH2 * 4 = C4H8